### v1

@@ Abstract
Quantum computing is regarded as a promising paradigm that may overcome the current computational power bottlenecks in the post-Moore era. The increasing maturity of quantum processors, especially superconducting ones, provides more possibilities for the development and implementation of quantum algorithms. As the crucial stages for quantum algorithm implementation, the logic circuit design and quantum compiling have also received significant attention, which covers key technologies such as quantum logic circuit synthesis (also widely known as quantum architecture search) and optimization, as well as qubit mapping and routing. Recent studies suggest that the scale and precision of related algorithms are steadily increasing, especially with the integration of artificial intelligence methods. In this survey, we systematically review and summarize a vast body of literature, exploring the feasibility of an integrated design and optimization scheme that spans from the algorithmic level to quantum hardware, combining the steps of logic circuit design and compilation optimization. Leveraging the exceptional cognitive and learning capabilities of AI algorithms, one can reduce manual design costs, enhance the precision and efficiency of execution, and facilitate the implementation and validation of the superiority of quantum algorithms on hardware.
keywords:
quantum circuit synthesis, quantum architecture search, quantum circuit optimization, quantum compiling, quantum machine learning, artificial intelligence
1 Introduction
@@
In recent years, quantum computing and quantum machine learning have garnered extensive attention. Due to their unique physical properties, quantum computing is considered the most promising new computing architecture to break through the existing computational power bottlenecks in the post-Moore era. Several notable quantum algorithms, such as Shor’s algorithm for large integer factorization [170] and Grover’s algorithm for unstructured search [64], have been theoretically proven to enjoy quantum superiority over classical algorithms. With the advent of the Noisy Intermediate-Scale Quantum (NISQ) era in the past decade [80, 157], quantum superiority has not only been demonstrated theoretically but also experimentally verified on quantum prototypes. For example, Google has demonstrated quantum superiority on a superconducting quantum [^1] For conciseness and clarity, when referring to quantum hardware in this article, the default is the superconducting system, which is a mainstream quantum hardware technology. prototype through random circuit sampling tasks [12, 197]. Similarly, the team with USTC has verified this through random boson sampling tasks on a photonic quantum prototype [211].
@@
Compared to random circuit sampling and random boson sampling, exploring the superiority of non-trivial and practically significant quantum algorithms presents a significant challenge, particularly given the constraints of existing quantum hardware, in terms of e.g., coherence time, noise, and physical topology. Many existing quantum algorithms utilize oracles [170, 64, 189], which are black-box unitary matrices where the corresponding unitary operation is known yet whose implementations are not explicitly defined. To best execute the algorithms, one needs to implement these unitary operators in the algorithms with minimal cost (such as circuit depth, gate count, etc.) and thus obtain the corresponding logic circuits of the algorithms. These logic circuits are further optimized with specific optimization goals. For instance, in the NISQ era, the focus is on optimizing two-qubit gates, while in the fault-tolerant quantum computing era, the emphasis shifts to optimizing T gates. Finally, the logic circuits must be compiled, requiring the optimal mapping from logical qubits to physical qubits and the addition of necessary SWAP gates to satisfy the circuit’s connectivity requirements. For variational quantum algorithms, we can also design and optimize parameterized quantum circuits, breaking traditional design paradigms. By integrating hardware constraints, we seek a balance between theoretical precision and noise (where reducing circuit depth decreases theoretical precision, but reducing noise can decrease execution errors) to achieve optimal results.
@@
Figure 1 shows the entire process from quantum algorithm design to execution. This process includes transforming a quantum algorithm into unitary transformations, generating logic circuits through synthesis and optimization methods, and then compiling these circuits, considering the physical qubit topology and other quantum hardware constraints, into executable quantum programs on target quantum processors. In other words, the transformation from a quantum algorithm to an executable quantum program can be divided into three steps: quantum algorithm design, quantum logic circuit design, and quantum compiling. In reviews of modern quantum computing technology [81], these steps are often collectively referred to as quantum circuit compilation, with relevant research briefly summarized. Early work also roughly summarized quantum circuit compilation in the NISQ era with heuristic methods [102], including genetic algorithms, genetic programming, ant colony optimization, and planning algorithms. In this paper, we focus on efficiently transforming quantum algorithms into executable quantum programs by comprehensively summarizing research on logic circuit synthesis and optimization, as well as qubit mapping and routing. Additionally, we provide prospects and insights into automatic quantum logic circuit design and compilation optimization, which enables researchers to focus more on algorithm design, fully leveraging the potential of NISQ-era quantum hardware and promoting the practical application and superiority verification of quantum algorithms.
@@
This survey first introduces and discusses existing major methods for representing quantum circuits. By representing circuits in various forms, such as gate model [46], directed acyclic graph [123], phase polynomial [7], ZX diagram [36], and tensor network [22], we can explore different synthesis, optimization, and compilation methods corresponding to these representation forms. For the problem of quantum logic circuit synthesis, we first examine the applications, which are divided into cases with and without target unitary matrices. Applications with given unitary matrices include oracle implementation [23, 158], while applications without target unitary matrices encompass variational quantum algorithms [63, 48, 195], error correction codes [33, 37], and nonlinear unit design [159]. Furthermore, we explore the methods for quantum circuit synthesis. We can generally classify the algorithms by dividing the task into two steps — deciding circuit structure and evolving rotation gate parameters (if applicable). The introduction of machine learning algorithms has made it possible to search for quantum circuits containing parameters, yielding better results than traditional methods.
For the problem of quantum logic circuit optimization, we have categorized relevant literature from the perspectives of optimization objectives and methods. As an important step following circuit synthesis, quantum circuit optimization aims to obtain a better equivalent circuit based on the given optimization objectives. Considering the differing characteristics of the NISQ era and the fault-tolerant quantum computing era, related studies have proposed various optimization objectives, which are also tied to the quantum circuit representation method. At the algorithm level, besides borrowing methods such as pattern matching and peephole optimization from traditional compilation optimization [127], we can also reduce the circuit optimization problem to other well-studied problems [75, 9, 166] or build an equivalent circuit database based on a specific set of quantum gates, obtaining the optimal subcircuit through searching the database [8, 99, 26]. Additionally, reinforcement learning methods have also recently achieved notable success in the field of quantum circuit optimization [56, 110, 161], and can even self-discover new matching patterns [163], surpassing traditional methods. We also summarize optimization methods for variational quantum algorithms. These studies, based on specific variational quantum algorithms [63, 177, 204, 117], propose corresponding optimization methods to make the optimized variational quantum circuits more suitable for quantum processors in the NISQ era.
In the quantum circuit compiling state, we need to consider the topological structure of the physical qubits, specifically the connectivity between them. Two-qubit quantum gates cannot be applied between qubits that lack a direct coupler, so some gates in the logic circuit obtained during the logical circuit design stage cannot be directly executed on a quantum computer. By adding SWAP gates to the circuit, one can exchange the information and state of two qubits; however, each SWAP gate requires three CNOT gates for implementation. Introducing too many SWAP gates into the circuit can severely impact operational accuracy. Therefore, we sorted out the algorithms for quantum qubit mapping and routing. Similar to logical circuit optimization, qubit mapping and routing focus on the number of quantum gates and circuit depth, but they also emphasize runtime and fidelity [178, 135, 144, 113, 114, 145]. Due to the optimization nature of the algorithm, qubit routing methods are akin to logical circuit optimization methods, including exact solutions [172], heuristic methods [217], reduction [135, 191, 130], and machine learning approaches [155, 171, 77].
@@
This leads to a key question explored in this paper: is it possible to develop an integrated algorithm that encompasses both quantum logical circuit design and quantum compiling? Such an algorithm should include the following steps: (1) Generate the optimal logical circuit corresponding to the quantum algorithm based on the built-in gate set of the quantum processor; (2) Complete the selection, mapping, and routing of quantum bits and couplers according to the calibration data of the quantum processor; (3) Balance theoretical accuracy and practical noise, achieving the optimal performance of the quantum processor with an appropriate trade-off in theoretical accuracy. Solving all these constraints is a challenging problem. However, the introduction of artificial intelligence may facilitate the design of related algorithms. Relevant literature indicates that with the introduction of machine learning algorithms, quantum logical circuit synthesis, optimization, and compilation have become significant application scenarios for AI4Science. Idealy, integrating all the above steps into the compiler of the quantum hardware cloud platform could automate these processes through AI algorithms. This would significantly enhance the practical performance of the submitted algorithms, reducing the additional labor costs post-quantum algorithm design, and providing a new paradigm for quantum computing applications.
@@
In summary, the contributions of this paper are as follows: (1) Conducting a comprehensive analysis of the various steps and challenges in the development and execution of quantum algorithms, covering the entire process from quantum logical circuit design to quantum compiling; (2) Systematically organizing the algorithms and their characteristics related to quantum logical circuit synthesis, optimization, as well as qubit mapping and routing; (3) Analyzing the requirements for real-world implementation of quantum algorithms in the NISQ era, and proposing an integrated quantum circuit design and compiling solution. This solution leverages artificial intelligence to enhance the performance and design efficiency of quantum algorithms and envisions a new paradigm for quantum compiling in the age of artificial intelligence.
Figure 2: Quantum Circuit Representation. Gate Model: most common representation method; DAG: example DAG transformed from the gate model example, “start” and “end” nodes are added to complete the graph; QMDD: QMDD representation of a three-qubit quantum operation, we present the transformation matrix and the layerwise structure as well as the size of the corresponding matrix of each node; Tensor Network: a quantum state |$\ket{\psi}$ can be represented by multiple common types of tensor network; ZX-Diagram: the green and red spider in ZX-diagram as well as an example ZX-diagram transformed from gate model example.
2 Quantum Circuit Representation
@@
We begin by discussing the representation of quantum circuits, which involves using graphs or vectors to depict specific information about the circuits. The representation of quantum circuits is crucial for their synthesis and optimization. Some synthesis and optimization algorithms are based on specific representations. Different representations of the same circuit can reduce the complexity of subsequent circuit design and compiling steps while still conveying the necessary information.
2.1 Quantum Gate Model
@@
Currently, the most common and fundamental method for representing quantum circuits is through the quantum gate model. This method characterizes quantum circuits using a system similar to classical gate circuits by representing the operations on qubits as quantum gates. A quantum gate can operate on multiple qubits. Each gate corresponds to a unitary transformation, modifying the quantum state and enabling circuit operations. Quantum gates can be divided into two main categories: parameterized and non-parameterized gates. The former allows for continuous transformations. To implement quantum algorithms, a set of quantum gates that can approximate any unitary matrix with arbitrary precision is required. This set is known as the universal quantum gate set [142]. According to [142], the necessary and sufficient condition for a universal quantum gate set is that it can approximate any single-qubit unitary transformation and includes a two-qubit gate. Other two or more qubit gates can be constructed from this universal set.
One of the most commonly used and well-studied universal quantum gate sets is Clifford+T [55]. The Clifford set includes H gates, S gates, and CNOT gates [60, 62], and quantum circuits containing only Clifford gates can be efficiently simulated by classical computers, making this set particularly important [61]. The Clifford set forms the foundation for many quantum algorithms and is defined as follows:
$\text{H}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix},\quad\text{S}=\begin{bmatrix}1&0\\ 0&e^{\text{i}\frac{\pi}{2}}\end{bmatrix}=\sqrt{\text{Z}},\quad\text{CNOT}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{bmatrix}.$ (1)
Through the combination of these three gates, we can generate a variety of other gates. For example, the Pauli-Z gate can be obtained by S2, the Pauli-X gate can be obtained by HZH, and the Pauli-Y gate can be obtained by SXS. However, these basic Clifford gates are not universal on their own. They cannot approximate many other gates with arbitrary precision within a finite sequence. To achieve universal quantum computation, we need to supplement the Clifford set with the T gate:
$\text{T}=\begin{bmatrix}1&0\\ 0&e^{\text{i}\frac{\pi}{4}}\end{bmatrix}=\sqrt{\rm S}=\sqrt[4]{\rm Z}.$ (2)
@@
The T gate is a phase gate. When we supplement the Clifford set with the non-Clifford T gate, the quantum gate set can continuously generate new non-Clifford gates, making the H, S, T, CNOT set capable of approximating any quantum gate. The Clifford+T set is the most commonly used universal quantum gate set, primarily because Clifford gates are relatively easy to implement on superconducting quantum computers. However, the T gates are more challenging to implement, so one major objective of quantum circuit optimization is to minimize the number of T gates required.
Besides the Clifford+T set, there are other general quantum gate sets, such as the parameterized rotation gates Rx$(\theta)$, Ry$(\theta)$, and Rz$(\theta)$, along with the CNOT gate, which is commonly used in quantum machine learning. The unitary matrices for the three parameterized rotation gates are as follows:
$\rm Rx(\theta)=\begin{bmatrix}\cos(\frac{\theta}{2})&-\text{i}\sin(\frac{\theta}{2})\\ -\text{i}\sin(\frac{\theta}{2})&\cos(\frac{\theta}{2})\end{bmatrix},\quad\rm Ry(\theta)=\begin{bmatrix}\cos(\frac{\theta}{2})&-\sin(\frac{\theta}{2})\\ \sin(\frac{\theta}{2})&\cos(\frac{\theta}{2})\end{bmatrix},\quad\rm Rz(\theta)=\begin{bmatrix}e^{-\text{i}\frac{\theta}{2}}&0\\ 0&e^{\text{i}\frac{\theta}{2}}.\end{bmatrix}$ (3)
Compared to the Clifford+T set, the main advantage of this universal gate set is that all single-qubit gates are parameterized, allowing for continuous transformations. As a result, the number of gates required to approximate arbitrary single-qubit gates is significantly reduced compared to using Clifford+T gates with discrete angles. However, the arbitrary angles of parameterized gates can affect accuracy when implemented on NISQ-era quantum computers. In the future fault-tolerant quantum computing era, it is also challenging to implement these gates at arbitrary angles through encoding.
Additionally, some universal quantum gate sets are designed for specific problems. For instance, the NCV quantum gate set is used to optimize circuits [129, 122, 124, 123], particularly multi-controlled-NOT gates. This set includes NOT gates, CNOT gates, and Controlled-V/V gates, where the unitary matrix for the V gate is
$\rm V=\frac{1+\text{i}}{2}\begin{bmatrix}1&-\text{i}\\ -\text{i}&1\end{bmatrix}.$ (4)
Some studies also use built-in gates on superconducting quantum computers directly, like the iSWAP$=e^{\text{i}\frac{\pi}{4}(XX+YY)}$ gate in [174]. These special cases will be explained in detail when encountered in the subsequent sections.
2.2 Directed Acyclic Graph

### v2

@@ Abstract
Quantum computing is a promising paradigm that may overcome the current computational power bottlenecks. The increasing maturity of quantum processors provides more possibilities for the development and implementation of quantum algorithms. As the crucial stages for quantum algorithm implementation, the logic circuit design and quantum compiling have also received significant attention, which covers key technologies, e.g., quantum logic circuit synthesis (also widely known as quantum architecture search) and optimization, as well as qubit mapping and routing. Recent studies suggest that the scale and precision of related algorithms are steadily increasing, especially with the integration of artificial intelligence methods. In this survey, we systematically review and summarize a vast body of literature, exploring the feasibility of an integrated design and optimization scheme that spans from the algorithmic level to quantum hardware, combining the steps of logic circuit design and compilation optimization. Leveraging the exceptional cognitive and learning capabilities of AI algorithms, it becomes more possible to reduce manual design costs, enhance the precision and efficiency of execution, and facilitate the implementation and validation of the superiority of quantum algorithms on hardware. An actively maintained paper list is available at: https://github.com/Thinklab-SJTU/awesome-ml4Qcircuit.
I Introduction
@@
In recent years, quantum computing and quantum machine learning have garnered extensive attention. Due to their unique physical properties, quantum computing is considered the most promising new computing architecture to break through the existing computational power bottlenecks in the post-Moore era. Several notable quantum algorithms, e.g. Shor’s algorithm for large integer factorization [1] and Grover’s algorithm for unstructured search [2], have been theoretically proven to enjoy quantum superiority over classical algorithms. With the advent of the Noisy Intermediate-Scale Quantum (NISQ) era in the past decade [3, 4], quantum superiority has not only been demonstrated theoretically but also experimentally verified on quantum prototypes. For example, Google has demonstrated quantum superiority on a superconducting quantum [^1] For clarity and conciseness, all references to quantum hardware in this article refer to superconducting systems, if not otherwise specified. prototype through Random Circuit Sampling (RCS) tasks [5, 6]. Similarly, the team with USTC has verified this through Gaussian Boson Sampling (GBS) on a photonic quantum prototype [7].
@@
Exploring the superiority of non-trivial and practically significant quantum algorithms beyond RCS and GBS presents a significant challenge, particularly given the constraints of current quantum hardware, including limited coherence time, noise, and physical topology. Many quantum algorithms rely on oracles [1, 2, 8], which are black-box unitary matrices with known operations but undefined implementations. Implementing these algorithms requires efficient realization of these unitary operators, minimizing key metrics such as circuit depth and gate count. The corresponding logic circuits are then optimized based on specific goals. In the NISQ era, the primary focus has been on optimizing two-qubit gates, while in the fault-tolerant quantum computing (FTQC) era, optimization targets, such as T gates, become more prominent. Once optimized, logic circuits must undergo compilation, which involves mapping logical qubits to physical qubits and inserting necessary SWAP gates to meet connectivity constraints. For variational quantum algorithms (VQAs), specific synthesis and optimization are needed for parameterized circuits. Balancing hardware constraints with algorithmic goals, we aim to optimize circuit performance by minimizing noise while maintaining theoretical precision, recognizing the trade-off between circuit depth and execution errors.
@@
Fig. 1 illustrates the entire process from quantum algorithm design to execution. This process includes transforming a quantum algorithm into unitary transformations, generating logic circuits through synthesis and optimization methods, and then compiling these circuits, considering the physical qubit topology and other quantum hardware constraints, into executable quantum programs on target quantum processors. In other words, the transformation from a quantum algorithm to an executable quantum program can be divided into three steps: quantum algorithm design, quantum logic circuit design, and quantum compiling. In reviews of modern quantum computing technology [9], these steps are often collectively referred to as quantum circuit compilation, with relevant research briefly summarized. Early work also roughly summarized quantum circuit compilation in the NISQ era with heuristic methods [10], including genetic algorithms, genetic programming, ant colony optimization, and planning algorithms. In this paper, we focus on efficiently transforming quantum algorithms into executable quantum programs by comprehensively summarizing research on logic circuit synthesis and optimization, as well as qubit mapping and routing. Additionally, we provide prospects and insights into automatic quantum logic circuit design and compilation optimization, which enables researchers to focus more on algorithm design, fully leveraging the potential of NISQ-era quantum hardware and promoting the practical application and superiority verification of quantum algorithms.
@@
This survey introduces methods for representing and optimizing quantum circuits, covering quantum circuit representations such as gate models [11], directed acyclic graphs [12], phase polynomials [13], ZX diagrams [14], and tensor networks [15]. We explore synthesis and optimization methods tailored to these representations. For quantum logic circuit synthesis, we discuss applications with and without target unitary matrices. Applications with given unitary matrices include oracle implementation [16, 17], while those without target unitary matrices include VQAs [18, 19, 20], error correction codes [21, 22], and nonlinear unit design [23, 24]. Quantum circuit synthesis typically involves determining circuit structure and evolving rotation gate parameters, with machine learning increasingly improving optimization over traditional methods. For quantum circuit optimization, we categorize relevant literature based on objectives and methods. Quantum circuit optimization improves circuits based on optimization goals and is influenced by the different characteristics of NISQ and FTQC eras. Methods include pattern matching, peephole optimization [25], and reductions to well-studied problems [26, 27, 28]. Recent advances in reinforcement learning [29, 30, 31] have further improved optimization by discovering new matching patterns, surpassing traditional approaches. Optimization methods for VQAs, such as those for specific VQAs [18, 32, 33, 34], enhance their suitability for NISQ-era quantum processors. In quantum circuit compilation, we address physical qubit connectivity and the need for SWAP gates to handle limited connectivity between qubits. While SWAP gates enable qubit state exchanges, they introduce overhead in terms of gate count and can impact operational accuracy. Therefore, quantum qubit mapping and routing algorithms are essential for optimizing gate counts, circuit depth, runtime, and fidelity [35, 36, 37, 38, 39, 40]. These methods, including exact solutions [41], heuristics [42], reductions [36, 43, 44], and machine learning approaches [45, 46, 47], focus on optimizing the performance of quantum circuits.
@@
This leads to a key question explored in this paper: is it possible to develop an integrated algorithm that encompasses both quantum logical circuit design and quantum compiling? Such an algorithm should include the following steps: (1) Generate the optimal logical circuit corresponding to the quantum algorithm based on the built-in gate set of the quantum processor; (2) Complete the selection, mapping, and routing of quantum bits and couplers according to the calibration data of the quantum processor; (3) Balance theoretical accuracy and practical noise, achieving the optimal performance of the quantum processor with an appropriate trade-off in theoretical accuracy. Solving all these constraints is a challenging problem. However, the introduction of artificial intelligence may facilitate the design of related algorithms. Relevant literature indicates that with the introduction of machine learning algorithms, quantum logical circuit synthesis, optimization, and compilation have become significant application scenarios for AI4Science. Ideally, integrating all the above steps into the compiler of the quantum hardware cloud platform could automate these processes through AI algorithms. Ideally, integrating these steps into a quantum hardware cloud platform’s compiler, powered by AI algorithms, could automate the process, significantly improving the practical performance of submitted algorithms and eliminating the need for manual intervention after algorithm design.
@@
In summary, the contributions of this paper are as follows: (1) Conducting a comprehensive analysis of the various steps and challenges in the development and execution of quantum algorithms, covering the entire process from quantum logical circuit design to quantum compiling; (2) Systematically organizing the algorithms and their characteristics related to all sub-tasks, providing abstract mathematical definitions for each step, facilitating the introduction of AI methods and lowering the entry barrier for newcomers; (3) Analyzing the requirements for real-world implementation of quantum algorithms in the NISQ era, and proposing an integrated quantum circuit design and compiling solution. This solution leverages artificial intelligence to enhance the performance and design efficiency of quantum algorithms and envisions a new paradigm for quantum compiling in the age of AI.
[…]
II Quantum Circuit Representation
@@
We begin by discussing the quantum circuit representations. This involves using graphs or vectors to depict specific circuit information, crucial for their synthesis and optimization. Different synthesis and optimization algorithms are designed based on particular representations. Choosing the right representation can reduce the complexity of subsequent design and compiling steps while still conveying the necessary information.
II-A Quantum Gate Model
@@
The most common and fundamental method for representing quantum circuits is through the quantum gate model. This method characterizes quantum circuits using a system similar to classical gate circuits by representing the operations on qubits as quantum gates. A quantum gate can operate on multiple qubits. Each gate corresponds to a unitary transformation, modifying the quantum state and enabling circuit operations. Quantum gates can be divided into two main categories: parameterized and non-parameterized gates. The former allows for continuous transformations. To implement quantum algorithms, a set of quantum gates that can approximate any unitary matrix with arbitrary precision is required. This set is known as the universal quantum gate set [48]. According to [48], the necessary and sufficient condition for a universal quantum gate set is that it can approximate any single-qubit unitary transformation and includes a two-qubit gate.
[…]
@@
Through the combination of these three gates, we can generate a variety of other gates. For example, the Pauli-Z gate can be obtained by S2, the Pauli-X gate can be obtained by HZH, and the Pauli-Y gate can be obtained by SXS. However, these basic Clifford gates are not universal on their own. They cannot approximate many other gates with arbitrary precision within a finite sequence. To achieve universal quantum computation, we need to supplement the Clifford set with the T gate: $\text{T}=\begin{bmatrix}1&0\\ 0&e^{\text{i}\frac{\pi}{4}}\end{bmatrix}=\sqrt{\rm S}=\sqrt[4]{\rm Z}$. The T gate is a phase gate. When we supplement the Clifford set with the non-Clifford T gate, the quantum gate set can continuously generate new non-Clifford gates, making the H, S, T, CNOT set capable of approximating any quantum gate. The Clifford+T set is the most commonly used universal quantum gate set, primarily because Clifford gates are relatively easy to implement on superconducting quantum computers. However, the T gates are more challenging to implement, so one major objective of quantum circuit optimization is to minimize the number of T gates required.
Besides Clifford+T, there are parameterized universal gate sets. Consider the parameterized rotation gates $\rm Rx(\theta)=\begin{bmatrix}\cos(\frac{\theta}{2})&-\text{i}\sin(\frac{\theta}{2})\\ -\text{i}\sin(\frac{\theta}{2})&\cos(\frac{\theta}{2})\end{bmatrix}$, $\rm Ry(\theta)=\begin{bmatrix}\cos(\frac{\theta}{2})&-\sin(\frac{\theta}{2})\\ \sin(\frac{\theta}{2})&\cos(\frac{\theta}{2})\end{bmatrix}$, and $\rm Rz(\theta)=\begin{bmatrix}e^{-\text{i}\frac{\theta}{2}}&0\\ 0&e^{\text{i}\frac{\theta}{2}}.\end{bmatrix}$. Along with the CNOT gate, they form a commonly used gate set in quantum machine learning. Compared to the Clifford+T set, the main advantage of this set is that all single-qubit gates are parameterized, allowing for continuous control over rotation angles. As a result, the number of gates required to approximate arbitrary single-qubit gates is significantly reduced compared to using Clifford+T gates with discrete angles. However, their physical implementation on hardware is limited by the resolution of the control electronics. This leads to inherent errors in gate operation, as the analog relationship between pulse duration and rotation angle cannot be realized with infinite precision.
[…]
II-B Directed Acyclic Graph